an astronomical telescope with small amplitude has an objective whose power is 2 diopters. the lens is placed 60 cm from the eye piece, when the telescope is adjusted for minimum eye strain. calculate the angular magnification of telescope.


1
Expert's answer
2021-09-24T09:25:43-0400

"\text{For lens:}"

"D_l = 2; f_o = \frac{1}{D}=0.5m"

"\text{For eyepiece}"

"f_e +f_o = h"

"h = 60cm=0.6m"

"f_e = 0.1m"

"\text{angular magnification }M"

"M= \frac{f_o}{f_e}= 5"

"\text{Answer :angular magnification }M= 5"